OpenStudy (anonymous):

x^2+7x-2=0 what are the solutions and what would be the x intercepts?

hero (hero):

\[x^2 + 7x = 2\] ....from there, complete the square

OpenStudy (anonymous):

would it be \[\sqrt{2},-\sqrt{2}\]

OpenStudy (dumbcow):

solutions and x-intercepts are the same thing... try to follow my steps from the last problem. take half of 7, square it, then add it to both sides ...

OpenStudy (anonymous):

but i add the o into the problem to get the intercepts right

OpenStudy (dumbcow):

??

OpenStudy (anonymous):

\[3.5^{2}\]

OpenStudy (dumbcow):

good

OpenStudy (anonymous):

\[7x+3.5^{2}=2+3.5^{2}\]

OpenStudy (dumbcow):

good, dont leave off the x^2 though Now factor left side

OpenStudy (anonymous):

ok id need to add x^2 first so it would be

OpenStudy (anonymous):

\[x ^{2}+7x+3.5\]

OpenStudy (dumbcow):

yes then it factors to \[(x+3.5)^{2}\]

OpenStudy (anonymous):

then i do the right side

OpenStudy (anonymous):

would that be 14.25

OpenStudy (dumbcow):

If you start with \[x^{2} + bx = c\] you will end up with \[(x + b/2)^{2} = c + (b/2)^{2}\]

OpenStudy (dumbcow):

yes 14.25, good

OpenStudy (anonymous):

now that is one solution or no

OpenStudy (dumbcow):

not yet, we have to get rid of the square how do we do that?

OpenStudy (anonymous):

do i now have to combine

OpenStudy (dumbcow):

you should have \[(x+3.5)^{2} = 14.25\] how do i get rid of the square

OpenStudy (anonymous):

square root of both sides then subtract?

OpenStudy (dumbcow):

perfect

OpenStudy (anonymous):

i can see how you want me to do it but im getting confused on how to accomplish it im sorry i must seem really stupid

OpenStudy (dumbcow):

perfect

OpenStudy (dumbcow):

whats confusing you? you told me how to do it...square root both sides then subtract then you will have x alone and you are done

OpenStudy (anonymous):

im looking at your example and trying to figure out if it should be in a fraction

OpenStudy (dumbcow):

perfect

OpenStudy (dumbcow):

sqrt(14.25) = +-3.775 x + 3.5 = +-3.775

OpenStudy (dumbcow):

perfect

OpenStudy (dumbcow):

oh i see, yes i used fractions all the way through and i came out nice but many times you have to use your calculator to find the square roots

OpenStudy (dumbcow):

perfect

OpenStudy (dumbcow):

so no you can write things as fraction or decimal, whatever you like \[x = \pm3.775 - 3.5\]

OpenStudy (anonymous):

so in the end im looking for the answer as a radical

hero (hero):

Fraction form is more exact

OpenStudy (anonymous):

the first part of my problem is asking for the solutions as exact anser using radicals rationalizing all denominators as nessecary express all complex tems in terms of i and using commas to separate answers as needed, then part two is to find the x intercepts it is showing it as a graphing problem

OpenStudy (dumbcow):

ok if you need an exact answer or in radical form then try to use fractions all the way through ie. instead of 3.5 use 7/2 3.5^2 = 49/4 our problem would be rewritten as: (x+7/2)^2 = 2 + 49/4 = 57/4 x + 7/2 = +-sqrt(57)/2 x = (+-sqrt(57) - 7)/2

OpenStudy (dumbcow):

remember the solution is the x-intercepts

OpenStudy (anonymous):

in previous ones i also had 0 in each is that still the case for this problem?